OpenStudy (anonymous):

I need to find the area of: \[D_5=\left\{ (x,y)\in \mathbb{R} : 0\le x \le x^2+y^2 \le 1\right\}\] Would it make sense to use polar coordinates? Like \[x=r*cos(a)\]and \[y=r*sin(a)\]?

OpenStudy (baru):

u mean area of a unit circle?

OpenStudy (anonymous):

Not sure, but since there is \[x^2+y^2\] It could be some kind of circle

OpenStudy (anonymous):

Usually I would just use Fobinis theorem, but I am having trouble because I have \[ 0\le x \le x^2+y^2 \le 1\] Instead of something where x and y are seperate

OpenStudy (baru):

i dont understand what sort of region we are looking at... but generally presence of \(x^2+y^2\) terms suggests that a switch to polar is convenient

OpenStudy (anonymous):

This is the region:\[D_5=\left\{ (x,y)\in \mathbb{R}^2 : 0\le x \le x^2+y^2 \le 1\right\}\]

OpenStudy (anonymous):

Yea, thats what I though aswell.

OpenStudy (anonymous):

So that would give: \[D_5=\left\{ (x,y)\in \mathbb{R} : 0\le r*cos(a) \le r^2 \le 1\right\}\] Now how would I make the double integral from this? - Using Fubinis theorem

OpenStudy (anonymous):

How do you know it is 0.5?

ganeshie8 (ganeshie8):

\(x \le x^2+y^2\) rearrange above inequality and put it in standard form

OpenStudy (anonymous):

Soo, to see the circles, you split it up like this: \[D_5=\left\{ (x,y)\in \mathbb{R} : 0\le x \le x^2+y^2 \le 1\right\}\]Becomes \[D_5=\left\{ (x,y)\in \mathbb{R} : 0\le x \le x^2+y^2,x \le x^2+y^2 \le 1\right\}\]

ganeshie8 (ganeshie8):

I guess yeah. Looks we need to include the region in IVth quadrant too

ganeshie8 (ganeshie8):

because \(0\le x\) gives you both I and IV quadrants

OpenStudy (anonymous):

Yea, so that would give an area of \[A(D_5)=\frac{\pi}{2}\]

ganeshie8 (ganeshie8):

shaded region = \(\frac{1}{2}[\pi*1^2 - \pi*(\frac{1}{2})^2] =\frac{3\pi}{8} \) right ?

OpenStudy (anonymous):

Yea, right \[A=\pi*r^2\],this cannot be the answer though. (It is an exam from last year, that I am practicing. And you are supposed to insert the missing number from 1-99) And I ahve \[A(D_5)=\frac{\pi}{insert}\]

ganeshie8 (ganeshie8):

that is so wrong, wait the center of smaller circle is not at origin

ganeshie8 (ganeshie8):

i mean, our earlier work is wrong

ganeshie8 (ganeshie8):

lets start over

OpenStudy (anonymous):

Aight, you dont think that I should be using Fubinis theorem here? First converting it into polar coordinates, then setting up the double integral?

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

before thinking of setting up an integral to find the area of region, we need to figure out exactly what area we need to find

ganeshie8 (ganeshie8):